OpenStudy (anonymous):

Let f(x) = x - 1 and g(x) = 2x + 3. What is f(g(x)) ?

OpenStudy (amistre64):

hmm, what would f(a) be?

OpenStudy (mertsj):

f(g(x)) means "replace each x in the function f with g(x)

OpenStudy (anonymous):

f(tiger) = tiger -1 f( 5 x)= 5 x -1 f(3x +4) = (3x+4) -1= 3 x+3 g(x)=2x +3 What is f(g(x))?